OpenStudy (anonymous):

You have three resistors in a parallel circuit with resistances 6 ohms, 12 ohms, and 24 ohms. What is the total resistance of the circuit?

OpenStudy (perl):

$$ \Large \frac{1}{R_{total}} =\frac{1}{R_{1}} + \frac{1}{R_{2}} +\frac{1}{R_{3}} $$

OpenStudy (perl):

so you can plug that in

OpenStudy (anonymous):

I did that, I plugged in 6, 12, and 24, but didn't know really what to do next

OpenStudy (perl):

$$\LARGE R_{total} =\frac{1}{\frac{1}{R_{1}} + \frac{1}{R_{2}} +\frac{1}{R_{3}}}$$

OpenStudy (anonymous):

but the answer is 24/7 and I have no idea how that worked out

OpenStudy (perl):

$$\LARGE R_{total} =\frac{1}{\frac{1}{6} + \frac{1}{12} +\frac{1}{24}}\\ \LARGE=\frac{1}{\frac{4}{24} + \frac{2}{24} +\frac{1}{24}}\\ \LARGE =\frac{1}{\frac{(4+2+1)}{24}} =\frac{1}{\frac{7}{24}} = \frac{24}{7} \\ $$

OpenStudy (perl):

$$\LARGE R_{total} =\frac{1}{\frac{1}{6} + \frac{1}{12} +\frac{1}{24}} \\[0.5 in] \text{use common denominator}\\[0.1in] \LARGE=\frac{1}{\frac{4}{24} + \frac{2}{24} +\frac{1}{24}}\\[0.1 in] \LARGE =\frac{1}{\frac{(4+2+1)}{24}} =\frac{1}{\frac{7}{24}} = \frac{24}{7} \\ $$